# Prompt-OT: An Optimal Transport Regularization Paradigm for Knowledge Preservation in Vision-Language Model Adaptation

### v1

@@ 3.3 Theoretical Analysis
Modeling cross-instance correlations can potentially characterize both within-class relations and between-class relations. Under the proposed OT-based constraint, each element in the transport map $\bm{\gamma}$ represents the amount of mass transported between two instances from the zero-shot representation and adapted representation. The learned transport map inherently captures cross-instance correlations because OT enforces a structured assignment, where similar instances are mapped with lower cost while maintaining the geometry of the data distribution. This process happens automatically because OT optimizes the movement of probability mass, ensuring that within-class relations remain compact while between-class relations reflect meaningful transitions, ultimately shaping a well-organized feature space.
Here, we focus more on how using OT-loss can potentially enlarge the feasible parameter space, which potentially results in easier optimization and better convergence.
We first give a general definition of the loss of point-wise constraint: {definition} Given $\bm{X}$ and $\bm{X}_{\texttt{zs}}$, the point-wise constraint is defined as:
@@
We observe that it can be presented in a similar form of optimal transport: {lemma} An equivalent form of Definition 3.3 is presented as:
@@
We first note that, clearly, according to Definition 3.2, $\hat{\gamma}\in\mathcal{U}(\bm{a},\bm{b})$ in Lemma 3.3 is a valid transport map. Accordingly,
@@
where $\mathcal{L}_{ce}$ denotes the cross-entropy loss widely used in adaptation for CLIP. $\mathcal{P}=\{\bm{P}^{l}\}_{l\in\mathcal{S}_{\texttt{vision}}},\mathcal{Q}=\{\bm{Q}^{l}\}_{l\in\mathcal{S}_{\texttt{text}}}$ denote all learnable tokens from different layers. $\lambda\geq 0$ is a tuning hyper-parameter.
This objective can be considered as the generalized Lagrange function of the following optimization problem:
$\min_{\mathcal{P},\mathcal{Q}}\quad\mathcal{L}_{ce}$ (14)
$s.t.~{}R(\bm{X},\bm{X}_{\texttt{zs}})-\epsilon^{2}\leq 0,$
@@
Theorem 3.3 immediately suggests that using optimal transport constraints allows a broader feasible parameter domain of learnable tokens $(\mathcal{P},\mathcal{Q})$. A larger feasible domain often indicates there are potential more local minimal existing and provides more pathways for the optimizer to explore, which increases the chances of finding a better solution and alleviates a heavy hyperparameter search, i.e., a uniform hyperparameter can achieve reasonable performance across different datasets.
@@ 3.4 Training and Inference
Table 1: Comparison with state-of-the-art methods on base-to-novel generalization. The best accuracies are bolded. HM indicates the harmonic mean. Prompt-OT represents the best accuracy achieved across different datasets by searching over epochs. It is not intended for direct comparison but is included here for reference.
@@ 4.1 Setups
We follow the experimental setups of [29, 9, 26] and evaluate on three tasks: Base-to-Novel Generalization, Cross-Data Evaluation, and Domain Generalization. For Base-to-Novel Generalization, we train on base classes and evaluate on both base and novel classes. Following [10], we use 11 datasets. To ensure fairness, we adopt two strategies: using the same epoch for all datasets (adhering to prior methods) and searching for the best epoch on each dataset (as a reference, only to show our method’s full potential). For Cross-Data Evaluation, we train on ImageNet [2] and test on 10 other datasets without extra fine-tuning. For Domain Generalization, we again train on ImageNet [2] and evaluate on four out-of-distribution datasets.
@@
Baselines. Despite compared with original CLIP [19], we select multiple recent prompt learning methods for comparison, including: CoOp [29], CoCoOp [28], ProDA [14], MaPLe [9], ProGrad [10], PromptSRC [10], PLOT [1], TCP [26], and QMaPLe/QCoOP [7]. It is worth mentioning that a portion of the methods are only designed for partial tasks, while our method does not have any obstacles to implementation for all tasks. For the sake of a fair comparison and proof of concept, we will skip comparing methods that introduce substantial additional knowledge (e.g., LLMs or larger CLIP models for knowledge distillation) in our main experiments, such as those proposed in [22, 13, 32, 24].
Implementation details. For fair comparison, We use a ViT-B/16-based CLIP model, consistent with prior work [9, 10, 26]. Prompts are initialized from a normal distribution, except for the text prompts in the text input, which begin with the embedding of ”a photo of”. We fix the learning rate at 0.005 and set $\lambda=10$ for all benchmarks. All reported results are the average of three runs, without any extra augmentation. Baseline results are taken from their respective papers. For more detail, refer to Appendix A.
@@ 4.2 Main Results
Our method demonstrates robust performance across three key evaluations. In the base-to-novel generalization setup, it substantially outperforms zero-shot CLIP and surpasses the previous state-of-the-art in base accuracy, novel accuracy, and harmonic mean. In cross-dataset evaluation, it achieves the highest average accuracy over 11 diverse datasets, again surpassing all baselines. Finally, in domain generalization, our method obtains the best overall accuracy without relying on prompt ensembles or text augmentation, underscoring both its effectiveness and simplicity.
@@ 5 Analysis
In this section, we will delve into a more detailed analysis of our proposed method. All experiments will be conducted consistently with the setups of base-to-novel generalization. For fairness, we exclude all other components, and only apply regularization on the IVLP.
Ours VS Point-wise Constraints. To illustrate the superiority of the proposed method over point-wise constraints, we evaluate it against several baselines: (i) naive $L_{2}$ constraints between $\bm{X}$ and $\bm{X}_{\texttt{zs}}$ with the same regularization strength $\lambda$, (ii) Adaptor-Cos [22], which introduces adapters at the end of the adapted encoders and employs cosine similarity as the point-wise constraint between representations, (iii) SRC [10], which applies $L_{1}$ losses for vision and text representations separately and a KL-divergence between prediction probabilities for each instance. (iv) Separate OT (see Appendix A), which applies OT-loss for vision and text representations separately. For a fair comparison, we exclude all additional augmentations, such as ensemble techniques and LLM-based augmentation. Our primary observation is that all these methods achieve a better balance between novel and base tasks than the model without any constraints, with a minimum improvement of 1.83% in harmonic mean accuracy. More importantly, the OT-based methods outperform all recent approaches based on point-wise constraints. Even without joint representation, applying OT for two domains separately can outperform all other point-wise methods. Additionally, compared to SRC, the previous state-of-the-art method, we achieved gains of 0.4%, 1.35%, and 0.93% in base task accuracy, novel task accuracy, and harmonic mean accuracy, respectively. Compared to $L_{2}$ constraints, we observe that with the same regularization strength $\lambda$, $L_{2}$ constraints, while preserving good generalizability to novel tasks, fail to yield any improvements on the base task. This suggests that $L_{2}$ constraints may be overly restrictive for model adaptation. In contrast, our proposed approach effectively balances performance across both the base and novel tasks, demonstrating that modeling cross-instance correlation while allowing the model greater flexibility in updating leads to an optimal trade-off, as expected in Fig. 1.
Table 4: Comparison of performance with point-wise constraints.
Reg. Method Base Novel HM
@@
Point-wise $L_{2}$ between $\bm{X},\bm{X}_{\texttt{zs}}$ 84.20 74.68 79.16
Adpator-Cos
From CoPormpt [22]
@@
Our 84.81 76.25 80.30
Effectiveness of Joint Representation. Here, we investigate alternative approaches to imposing constraints between zero-shot and adapted representations. Specifically, we examine four different strategies: (i) imposing the constraint solely on the text representation, (ii) imposing the constraint solely on the vision representation, (iii) applying constraints separately to text and vision representation distributions (i.e., Separate OT), and (iv) our proposed joint representation, which considers an instance includes both vision and text representations. As shown in Table 5, applying constraints to both modalities yields a significantly greater improvement (2.58% and 3.14% in HM) compared to imposing the constraint on a single modality (0.62% and 0.96% in HM). Moreover, our proposed joint representation obtain more gain than applying constraints on vision and text distributions separately. Our method leads to the highest increase in HM (3.14%), suggesting that it provides the most effective balance between minimizing errors in both novel and base classes.
@@
Sensitivity of $\lambda$. We further conduct a study to evaluate the impact of different values of $\lambda$ on adaptation. As shown in Fig. 3, as $\lambda$ increases, the constraints become stronger, leading to a decline in adaptation to the base class. Meanwhile, the accuracy on the novel class initially increases before decreasing, indicating that the zero-shot weights may not be optimal for both novel and base tasks. This suggests that some relaxation is necessary to achieve a well-balanced trade-off. Excessively strong constraints may prevent the model from reaching the low-error regions of both tasks in the optimization landscape. Empirically, we find that $\lambda=10$ achieves a good balance between generalization and adaptation.
@@ Appendix A Implementation Detail
We train PromptOT for 30 epochs for the Base-to-Novel Generalization benchmark and 2 epochs for the remaining three benchmark settings, respectively. The respective epochs are fixed across all datasets.
Following a deep prompting strategy, we apply prompts to all transformer blocks for the first benchmark and the first three transformer blocks for the remaining benchmarks, and we use four text prompts and four vision prompts.
Following previous works [10], in For Base-to-Novel Generalization benchmark, we set the first half of classes as base classes and the remaining classes as novel classes. Models are only trained on base classes. The accuracy is reported based on the evaluation of base classes and novel classes, respectively.
A.1 Separate OT

### v2

@@ 3.3 Theoretical Analysis
Modeling cross-instance correlations can potentially characterize both within-class relations and between-class relations. Under the proposed OT-based constraint, each element in the transport map $\bm{\gamma}$ represents the amount of mass transported between two instances from the zero-shot representation and adapted representation. The learned transport map inherently captures cross-instance correlations because OT enforces a structured assignment, where similar instances are mapped with lower cost while maintaining the geometry of the data distribution. This process happens automatically because OT optimizes the movement of probability mass, ensuring that within-class relations remain compact while between-class relations reflect meaningful transitions, ultimately shaping a well-organized feature space. Here, we focus more on how using OT-loss can potentially enlarge the feasible parameter space, which potentially results in easier optimization and better convergence. We first give a general definition of the loss of point-wise constraint. Given $\bm{X}$ and $\bm{X}_{\texttt{zs}}$, the point-wise constraint is defined as:
@@
We observe that it can be presented in a similar form of optimal transport:
@@
We first note that, clearly, according to Eq. 5, $\hat{\gamma}\in\mathcal{U}(\bm{a},\bm{b})$ in Eq. 7 is a valid transport map. Accordingly,
@@
where $\mathcal{L}_{ce}$ denotes the cross-entropy loss widely used in adaptation for CLIP. $\mathcal{P}=\{\bm{P}^{l}\}_{l\in\mathcal{S}_{\texttt{vision}}},\mathcal{Q}=\{\bm{Q}^{l}\}_{l\in\mathcal{S}_{\texttt{text}}}$ denote all learnable tokens from different layers. $\lambda\geq 0$ is a tuning hyperparameter. This objective can be considered as the generalized Lagrange function of the following optimization problem:
$\min_{\mathcal{P},\mathcal{Q}}\quad\mathcal{L}_{ce}\quad s.t.~R(\bm{X},\bm{X}_{\texttt{zs}})-\epsilon^{2}\leq 0,$ (10)
@@
Theorem 3.3 immediately suggests that using optimal transport constraints allows a broader feasible parameter domain of learnable tokens $(\mathcal{P},\mathcal{Q})$. A larger feasible domain often indicates there are potentially more local minima existing and provides more pathways for the optimizer to explore, which increases the chances of finding a better solution and alleviates a heavy hyperparameter search, i.e., a uniform hyperparameter can achieve reasonable performance across different datasets.
@@ 3.4 Training and Inference
Table 1: Comparison with state-of-the-art methods on base-to-novel generalization. The best accuracies are bolded. HM indicates the harmonic mean. Prompt-OT represents the best accuracy achieved across different datasets by searching over epochs. It is not intended for direct comparison but is included here for reference. Wilcoxon signed-rank test is shown in Appendix B.
@@ 4.1 Setups
We follow the experimental setups of [44, 16, 41] and evaluate on three tasks: Base-to-Novel Generalization, Cross-Data Evaluation, and Domain Generalization. For Base-to-Novel Generalization, we train on base classes and evaluate on both base and novel classes. Following [17], we use 11 datasets. To ensure fairness, we adopt two strategies: using the same epoch for all datasets (adhering to prior methods) and searching for the best epoch on each dataset (as a reference, only to show our method’s full potential). For Cross-Data Evaluation, we train on ImageNet [5] and test on 10 other datasets without extra fine-tuning. For Domain Generalization, we again train on ImageNet [5] and evaluate on four out-of-distribution datasets. Please refer to Appendix A for more details.
@@
Baselines. Despite compared with original CLIP [30], we select multiple recent prompt learning methods for comparison, including: CoOp [44], CoCoOp [43], ProDA [22], MaPLe [16], ProGrad [17], PromptSRC [17], PLOT [2], TCP [41], and QMaPLe/QCoOP [11]. It is worth mentioning that a portion of the methods are only designed for partial tasks, while our method does not have any obstacles to implementation for all tasks. For the sake of a fair comparison and proof of concept, we will skip comparing methods that introduce substantial additional knowledge (e.g., LLMs or extra larger CLIP models for knowledge distillation) in our main experiments, such as those proposed in [34, 21, 46, 38].
Implementation details. For fair comparison, we use a ViT-B/16-based CLIP model, consistent with prior work [16, 17, 41]. Prompts are initialized from a normal distribution, except for the text prompts in the text input, which begin with the embedding of ”a photo of”. We fix the learning rate at 0.005 and set $\lambda=10$ for all benchmarks. All reported results are the average of three runs, without any extra augmentation. Baseline results are taken from their respective papers. For more details, refer to Appendix A.
@@ 4.2 Main Results
Overall, our method demonstrates robust performance across three key evaluations. In the base-to-novel generalization setup, it substantially outperforms zero-shot CLIP and surpasses the previous state-of-the-art in base accuracy, novel accuracy, and harmonic mean. In cross-dataset evaluation and in domain generalization, our method obtains the best overall accuracy without relying on prompt ensembles or text augmentation, underscoring both its effectiveness and simplicity.
@@ 5 Analysis
In this section, we will delve into a more detailed analysis of our proposed method. All experiments will be conducted consistently with the setups of base-to-novel generalization. For fairness, we exclude all other components and only apply regularization on the IVLP.
Ours VS Point-wise Constraints.
To illustrate the superiority of the proposed method over point-wise constraints, we evaluate it against several baselines: (i) naive $L_{2}$ constraints between $\bm{X}$ and $\bm{X}_{\texttt{zs}}$ with the same regularization strength $\lambda$, (ii) Adaptor-Cos [34], which introduces adapters at the end of the adapted encoders and employs cosine similarity as the point-wise constraint between representations, (iii) SRC [17], which applies $L_{1}$ losses for vision and text representations separately and a KL-divergence between prediction probabilities for each instance.
Table 4: Comparison of performance with point-wise constraints. Reported $p$-values are obtained from Wilcoxon signed-rank tests comparing the baseline and our method.
Reg. Method Base Novel HM $p<0.05$
@@
Point-wise
$L_{2}$ between $\bm{X},\bm{X}_{\texttt{zs}}$ 84.20 74.68 79.16
Adaptor-Cos
From CoPrompt [34]
@@
Ours 84.81 76.25 80.30
(iv) Separate OT (see Appendix A), which applies OT-loss for vision and text representations separately. For a fair comparison, we exclude all additional augmentations, such as ensemble techniques and LLM-based augmentation. Table 4 reports both the performance metrics and the corresponding statistical test results. Our primary observation is that all these methods achieve a better balance between novel and base tasks than the model without any constraints, with a minimum improvement of 1.83% in harmonic mean accuracy. More importantly, the OT-based methods outperform all recent approaches based on point-wise constraints. Even without joint representation, applying OT for two domains separately can outperform all other point-wise methods. Additionally, compared to SRC, the previous state-of-the-art method, our proposed method achieved statistically significant gains of 0.4%, 1.35%, and 0.93% in base task accuracy, novel task accuracy, and harmonic mean accuracy, respectively. Compared to $L_{2}$ constraints, we observe that with the same regularization strength $\lambda$, $L_{2}$ constraints, while preserving good generalizability to novel tasks, fail to yield any improvements on the base task. This suggests that $L_{2}$ constraints may be overly restrictive for model adaptation. In contrast, our proposed approach effectively balances performance across both the base and novel tasks, demonstrating that modeling cross-instance correlation while allowing the model greater flexibility in updating leads to an optimal trade-off, as expected in Fig. 1.
Effectiveness of Joint Representation.
Here, we investigate alternative approaches to imposing constraints between zero-shot and adapted representations. Specifically, we examine four different strategies: (i) imposing the constraint solely on the text representation, (ii) imposing the constraint solely on the vision representation, (iii) applying constraints separately to text and vision representation distributions (i.e., Separate OT), and (iv) our proposed joint representation, which considers an instance includes both vision and text representations.
@@
As shown in Table 5, applying constraints to both modalities yields a significantly greater improvement (2.58% and 3.14% in HM) compared to imposing the constraint on a single modality (0.62% and 0.96% in HM). Moreover, our proposed joint representation yields greater improvements compared to applying constraints separately on vision and text distributions, which potentially results in unbalanced constraint strength. Our method leads to the highest increase in HM (3.14%), suggesting that it provides the most effective balance between minimizing errors in both novel and base classes.
@@
Sensitivity of $\lambda$.
We further conduct a study to evaluate the impact of different values of $\lambda$ on adaptation. As shown in Fig. 2, as $\lambda$ increases, the constraints become stronger, leading to a decline in adaptation to the base class. Meanwhile, the accuracy on the novel class initially increases before decreasing, indicating that the zero-shot weights may not be optimal for both novel and base tasks. This suggests that some relaxation is necessary to achieve a well-balanced trade-off. Excessively strong constraints may prevent the model from reaching the low-error regions of both tasks in the optimization landscape. Empirically, we find that $\lambda=10$ achieves a good balance between generalization and adaptation.
Efficiency.
Our method introduces only a minor computational overhead during training, as shown in Table 6, incurring approximately a 1.8% increase in runtime with negligible additional GPU usage.
Table 6: Comparison of Computational Cost.
Runtime GPU
w/o Constraint 0.0385s/step 3948MB
Ours 0.0392s/step 3950MB
@@ Appendix A Implementation Detail
We train PromptOT for 30 epochs for the Base-to-Novel Generalization benchmark and 2 epochs for the remaining three benchmark settings, respectively. The respective epochs are fixed across all datasets. All experiments are run on a node of the cluster with one V100 16GB NVIDIA GPU. Following a deep prompting strategy, we apply prompts to all transformer blocks for the first benchmark and the first three transformer blocks for the remaining benchmarks, and we use four text prompts and four vision prompts. Following previous works [17], in For Base-to-Novel Generalization benchmark, we set the first half of classes as base classes and the remaining classes as novel classes. Models are only trained on base classes. The accuracy is reported based on the evaluation of base classes and novel classes, respectively. The overview of this framework is shown in Fig. S3.
Figure S3: The overview of our proposed framework. Only the prompt tokens are trainable, and the rest of the weights in both the zero-shot encoders and the adapted encoders are frozen. Despite the cross-entropy $\mathcal{L}_{ce}$ adopted, we also adopt the proposed joint optimal transport loss $\mathcal{L}_{\mathrm{jot}}$ between joint zero-short representation and adapted representation to constrain the model.
A.1 Datasets
We used the following datasets in our experiments:
Exps. 1 and 2. ImageNet [5], Caltech101 [7], OxfordPets [29], StanfordCars [20], Flowers102 [28], Food101 [1], and FGVCAircraft [24], SUN397 [39], UCF101 [35], DTD [3], and EuroSAT [12].
Exp. 3. ImageNet [5] as a source dataset and use ImageNet-A [14], ImageNet-R [13], ImageNet-Sketch [37] and ImageNetV2 [32].
A.2 Separate OT
@@ A.3 Reproducibility
Appendix B Wilcoxon signed-rank test
We perform a Wilcoxon signed-rank test [4] to compare our method with PromptSRC, the previous state-of-the-art, using harmonic mean accuracy as the evaluation metric. As shown in Table S7, our method achieves statistically significant improvements over PromptSRC (p-value ¡ 0.05).
Table S7: Wilcoxon signed-rank test on Exp 1.
PromptSRC VS Prompt-OT
P-value 0.016
Appendix C Limitation
While our proposed PromptOT framework demonstrates strong performance across various benchmarks and introduces theoretically grounded improvements over prior prompt learning approaches, we acknowledge certain limitations that open avenues for future exploration:
First, the OT-based regularization introduces additional computational overhead compared to simpler constraints (e.g., point-wise L2 losses). Although we adopt mini-batch OT solvers to keep training efficient, the method still requires access to additional computational resources.
Second, while our approach avoids relying on external augmentations or auxiliary models (e.g., large LLMs), this also means it may not fully leverage certain recent advances in generative or data-enhancement techniques. Integrating such components in a resource-efficient manner could further boost performance.